OpenStudy (anonymous):

solve dis one pls 3^lnx=2^lny if x' nd y' r solution for it den wats x'

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

x= y^(ln2/ln3)

OpenStudy (anonymous):

3^lnx = 2^lny take log to base 3 both side then lnx = log(base3)(2^lny) x = e^( log(base3)(2^lny)) x = 2^(lny/ln3)

OpenStudy (anonymous):

i didnt get u bro

OpenStudy (anonymous):

were do u have difficulty

OpenStudy (anonymous):

i hav doubt rite frm first step its 3 power lnx.....

OpenStudy (anonymous):

as i mentioned the steps involved are 1) take log to base 3 of both sides 2)u get lnx now for example consider ln(r) = t ,then r = exponential(t) i.e, r=e^t 3)then u have x = e^(log(base 3) (2^lny)) now as per property of logarithms log(base x)y = lny/lnx or in general log(base x)y = (log(base z)y)/(log(base z)x) so u get x =e^( (ln(2^lny)) /ln(3)) so this can be solved and written as x = 2^(ln(y) / ln(3)) i hope u got it.